OpenStudy (anonymous):

An electron is accelerated from a very low velocity )0 speed) by the application of potential difference of V volts. if the de broglie wavelenght should change (ie decrease) by 1 % what % increase in V cause it?

OpenStudy (ujjwal):

If i remember it correctly, we had a relation like this: \[\lambda=\frac{12.27 A^0}{\sqrt V}\]where \(\lambda\) is De-broglie wavelength and V is Potential difference.

OpenStudy (anonymous):

man...wat is this...!!

OpenStudy (ujjwal):

\(A^0\) is angstrom equivalent to \(10^{-10}\)

OpenStudy (anonymous):

Did u get my question @ujjwal

OpenStudy (ujjwal):

Dude, I gave you the relation. Apply it! Decrease \(\lambda\) by 1% and see what % increase in V does it bring..

OpenStudy (ujjwal):

And yeah, that relation is exclusively for electrons.. You can use this relation in all cases..\[\lambda=\frac{h}{\sqrt{2mqV}}\]

OpenStudy (anonymous):

it is said that velocity is 0

OpenStudy (ujjwal):

Read the question 10 times and see what is zero...

OpenStudy (anonymous):



OpenStudy (ujjwal):

Have you seen this relation? \[\lambda=\frac{h}{P}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but the queestion tells abt energy

OpenStudy (ujjwal):

And \[P=\sqrt{2mKE}\]And \[KE=qV\]So, \[P=\sqrt{2mqV}\]So, \[\lambda=\frac{h}{\sqrt{2mqV}}\]

OpenStudy (anonymous):

yhxx